Form the cross product of vector A = 23km 37° and vector C = 47km 193°


1
Expert's answer
2021-03-29T09:01:45-0400

Let's first find the "x"- and "y"-components of vectors A and C:


"A_x=Acos\theta=23\ km\cdot cos37^{\circ}=18.37\ km,""A_y=Asin\theta=23\ km\cdot sin37^{\circ}=13.84\ km,""C_x=Ccos\theta=47\ km\cdot cos193^{\circ}=-45.8\ km,""C_y=Csin\theta=47\ km\cdot sin193^{\circ}=-10.6\ km."

We can find the cross product of vector A and vector C as follows:


"\vec{A}\times \vec{C}=\begin{vmatrix}
 A_x & A_y \\
 C_x & C_y
\end{vmatrix}\cdot\vec{k},""\vec{A}\times \vec{C}=(A_xC_y-C_xA_y)\cdot\vec{k},""\vec{A}\times \vec{C}=(18.37\ km\cdot(-10.6\ km)-(-45.8\ km)\cdot13.84\ km)\cdot\vec{k},""\vec{A}\times \vec{C}=(439.15\ km^2)\cdot\vec{k}"